OpenStudy (anonymous):

if a pair of fair dice is rolled, what are the odds in favor of the sum of the numbers showing being 5?

Parth (parthkohli):

With each of the numbers, we have 6 combos. So, the total outcomes are 6 * 6 = 36. We can have a sum of 5 in these cases: {4,1}{3,2}{2,3}{1,4} --> 4 cases. Probability = \(\Large \text{Preferred number of outcomes} \over \Large \text{Total number of outcomes } \)

Parth (parthkohli):

That is ---> \(\Large \color{midnightblue}{\rightarrow {4 \over 36} = {1 \over 9}}\)

OpenStudy (anonymous):

so 1/9 are the odds in favor

Parth (parthkohli):

Yep.

Parth (parthkohli):

I mean the odds in favour are {4,1};{3,2};{2,3};{1.4}

OpenStudy (anonymous):

ty partrhkohli

OpenStudy (anonymous):

plz correct me.. if im not wrong then odds in favours are:{2,3}{3,2}...? does odds on favour means sum of two odd numbers...?

Parth (parthkohli):

@Mamoona_akbar no, odds in favour in this case means that what are the cases in which we get our required outcome, which in this case is a sum of 5.

OpenStudy (mertsj):

odds = favorable over unfavorable so 4/32=1/8

OpenStudy (anonymous):

im glad im not the only one not quite sure but parths answer 4 1 3 2 2 3 1 4 seems right

OpenStudy (anonymous):

Mertsj :i couldn't understood ur answer..

OpenStudy (mertsj):

What is the definition of odds?

OpenStudy (mertsj):

That is the key to understanding my answer.

OpenStudy (anonymous):

favorable over unfavorable

OpenStudy (anonymous):

i believe what are the odds if the dice are rolled that 5 will be the total

OpenStudy (anonymous):

im thinking 1/9 is the answer

OpenStudy (anonymous):

how do i post another question

OpenStudy (anonymous):

@Mertsj explain plz...

OpenStudy (anonymous):

as parth put it 6 combos on each die so 6*6=36 with 4 possible outcomes equalling 5 4 1 3 2 2 3 1 4 so 4/36=1/9

OpenStudy (mertsj):